An object is released from a height of 72 feet and is given an initial velocity of 22 ft/s.
When the object strikes the ground, how far has it traveled in the horizontal direction?
1
Expert's answer
2020-10-12T07:50:53-0400

The distance travelled by the object in the vertical direction is:


"h = \dfrac{gt^2}{2}"

If the height is "h = 72\space feet," and "g = 29.5\space ft/s^2" is the gravitational acceleration, then the falling time will be:


"t_{fall} = \sqrt{\dfrac{2h}{g}}"

Object's horizontal speed has not been changed with time. Thus, the distance in the horizontal direction travelled in time of falling will be:


"d = v_0t_{fall} = v_0\sqrt{\dfrac{2h}{g}}"

where "v_0 = 22ft\/s". Thus, obtain:


"d = v_0\sqrt{\dfrac{2h}{g}} = 22\cdot \sqrt{\dfrac{2\cdot 72}{29.5}} \approx 48.6\space feet"

Answer. 48.6 feet.